# Design of a Quality Management System based on the EU Artificial Intelligence Act

### v1

@@ 3.3 Prototype QMS - Version 1
Sub-Service 01: Risk Management System. As mentioned, the main focus of the prototype QMS is on the RMS sub-service. The overall structure of the RMS is based on the ISO 31000 standard [15] which is also applied to the AI system RMS presented in [33]. It follows a “Plan-Do-Act-Check” principle: Plan involves planning the RMS based on the specific use case and defining the processes’ sub-activities, Do encompasses conducting the designed RMS process, Act entails inspecting and evaluating the entire RMS process, and output, and Check involves refining and optimizing the RMS process. Additionally, [26] establishes five design principles for RMS for AI systems, which should also be integrated. These principles include i) incorporating multi-perspective expert assessment, such as drawing insights from various domains and involving AI experts in the process, ii) encouraging participation from diverse stakeholders in the risk assessment process, iii) identifying risks based on real-life scenarios, and iv) analyzing risks using metrics beyond accuracy, v) ensuring “human in the loop” processes for black-box models. This RMS can be seen as a prototype within the prototype QMS. The construction process of this RMS, according to ISO 31000, consists of an iterative process containing six sub-activities: component selection, risk identification, verification data selection, risk analysis, risk assessment, and risk mitigation. In the component selection, the users should be encouraged to select the AI system (in this case, the LLM) that is analyzed, and the task the AI system performs. In the second step, the user conducts a risk identification based on the risk categorization strategy for AI systems presented by [12]. A vocabulary-based approach is used where the user adds values for the LLM’s domain, purpose, capabilities, LLM user, and LLM subject. Based on these selected values, the risk class of the AI system is determined. A basic algorithm was implemented to evaluate the risk class according to the categorization presented by [12]. Future work will explore changing the risk identification process to a more stakeholder-oriented approach, as mentioned in design principle i) by [26]. Additionally, it will be considered whether risk identification should be designed to identify foreseeable risks and misuses, assuming the AI system is already classified as high-risk, instead of categorizing the AI system into a risk class. In the third step, the user can quantitatively analyze the AI system’s performance, explainability, and consistency risks.
Figure 3: QMS - RMS - Analysis
As shown in the UI screenshot in Fig. 3, the user can choose between different technical evaluation metrics applied in the risk analysis to the selected LLM. For performance metrics, the user can choose: i) Accuracy score (cf. [2]), ii) Rouge-n score (cf. [21]), and iii) Perplexity (cf. [14]). For explainability, a gradient-based saliency-map metric can be employed (cf. [20], initially invented for image classifiers: [30]). For consistency, a gradient-based adversarial example input modification method can be selected (cf. [34], initially invented for image classifiers: [32, 13]). In the fourth step, the model calculates a risk assessment documentation based on the risk identification and analysis results. This documentation aims to verify compliance with EU AIA regulations. The current documentation will not serve as proof, but the basic principle and the feasibility of automatically creating and downloading a risk and technical documentation from the prototype QMS can be demonstrated. In the final step, the user has the option to add strategies to mitigate assessed risks. This feature still needs to be fully developed and integrated into the prototype’s second version.
@@
To provide an example of how such a risk analysis and assessment result looks, the process was demonstrated with test verification data, and all presented metrics were selected to be computed. As LLM the Phi-3-mini-128k-instruct  [^18] Phi-3-mini: https://huggingface.co/microsoft/Phi-3-mini-128k-instruct, accessed on 30 July 2024 from Microsoft containing between 4 and 5 billion parameters was analyzed on the domain “Industry Process Description”, and the task “Summarization”. The model input was: “Extract the Actor and Activity pairs from the text. Return only the list of JSON documents in the following format: [’actor’: ’example_actor_1’, ’activity’: ’example_activity_1’, ’actor’: ’example_actor_2’, ’activity’: ’example_activity_2’, …] without any further explanation: The user creates a new process instance, then the system can execute the instance and stop it afterward.”.
@@
In Fig. 4, the performance results of the model from the risk assessment UI page are shown. The performance results display the numerical values for the model’s accuracy, Rouge score, and perplexity for the given input tasks. In Fig. 5, the explainability result is visualized as a saliency map, where each input token is colored between red (unimportant, less sensitive) and green (important, sensitive) to the model’s generated output.
@@
Sub-Service 02: Data Management and Governance System. The DMDGS is the second sub-service in the QMS. In the current first version of the prototype QMS, the DMDGS is kept very basic. It consists of two components in the UI: the data check page and the past data check page. The former, as shown in Fig. 7, allows the user to select the AI system (in this case, LLM) for which training, validation, or testing data has been used. Furthermore, the user can specify the name of the dataset, the origin, the type, the domain, and the size. To provide proof that the data is checked and compliant under Article 10 EU AIA, the user has a text field in which he/ she can enter a textual reference to the proof.
Figure 7: QMS - DMDGS - Data Check
In future work, the goal is to enhance the Data Management and Governance System (DMDGS) to provide functionality for analyzing and assessing datasets for training, validation, and testing quantitatively within the QMS. This includes evaluating datasets for biases and errors. Additionally, the QMS should verify that dataset splits are well-determined and that the data aligns with the goals of AI system development. It is also essential to check whether other EU data regulations, such as the GDPR [^19] GDPR - Website: https://gdpr-info.eu, accessed on 30 July 2024, need to be considered within the QMS.
@@ 4 Evaluation
The implementation and realization of functional (FR) and non-functional requirements (NFR) in the QMS are evaluated technically through a prototypical implementation in Sect. 3.3. In the sequel, FRs are evaluated based on user story scenarios and NFRs are qualitatively assessed based on technical evaluations and calculations regarding the required GPU storage and performance for the LLMs used.
FR Evaluation – User Stories:. A user story has the form: “As a [provider, deployer, user…] [I want to…], [ensure, guarantee, provide…]…”. LR01LR10 and SDR01SDR04 are evaluated along with ten user story scenarios that exemplify potential stakeholders’ needs within the QMS. The user stories are mainly structured to evaluate the constructed requirements chronologically. The evaluation also includes potential limitations, unimplemented requirements, and future work.
1As a provider or deployer of a high-risk AI or GPAI system, I want a QMS in the form of a web application where I can directly perform technical evaluations on the AI system (model), document these evaluations, and include data references and checks for used train, validate and test datasets of the AI system (model).” (LR01, LR02). A prototype has been developed to provide an initial version of this QMS as a web application. While the prototype meets the requirements specified in the user story, several optimizations are needed. For example, these include improving the risk identification process, enabling simultaneous participation from multiple stakeholders in the RMS, and incorporating technical checks on datasets used for AI (training) development.
2As an AI Expert, I want to compute, download, and assess the technical evaluation results of the AI system in charge, discuss with legal experts the sufficiency of the model results, and see the current results in a certain domain or task for potential improvement.” (LR02, LR08, LR09). The latter requirement mandates performing several technical evaluations on the AI system. For this QMS, LLMs have been used as an example of AI systems because they are already pre-trained and freely available on Hugging Face, making it easy to access and integrate them into the prototype QMS. Five different metrics have been implemented: three to evaluate the model’s performance, one to evaluate its explainability, and one for its consistency (as a general term for robustness and reliability). However, no standard protocols such as the ISO 4213 [17] for AI performance evaluation, ISO CD TS 6254 [18] for AI explainability evaluation, and ISO TR 24029 for AI robustness evaluation [16] have been used as a guideline. The metrics and calculations are currently included in the risk analysis component. However, a different sub-service can also be created for LR08 (Article 15 EU AIA) to prove the AI system’s accuracy, robustness, and cybersecurity separately. For that, the already implemented verification component can be reused. The download feature to comply with LR09 (Article 11 EU AIA) already exists. However, implementing requirements outlined by AI and legal experts to define the scope and design of such technical documentation would further enhance its quality.
3As a provider or deployer of a high-risk AI or GPAI system, I want at least one AI expert from the development team, a domain expert, and a legal expert to perform a risk assessment together in a certain domain for specific tasks, to identify foreseeable risks or misuses and to effectively mitigate the risk” (LR03). An RMS is implemented using the ISO 31000 risk management standard [15] to guide and structure the process. The implementation includes essential risk identification, risk analysis with technical evaluation metrics on the LLM, and a risk assessment page. However, improvements are needed in the risk identification process, a solution to incorporate multiple stakeholders must be established, and the mitigation strategy page must still be implemented.
4As an AI expert of the development team, I want to have a sub-service included in the QMS where I can reference and check the data used for the AI system’s development, or include a data check, to ensure that my training, validation, and testing data is compliant with all privacy and data ownership regulations defined in legal frameworks. This will provide transparency and proof that can be shared with legal authorities and internally with all AI development members.” (LR04). The QMS includes the presented DMDGS, but the user can only add textual descriptions of the data and the corresponding checks. Further improvements are required, such as uploading certified data checks and more in-depth analysis capabilities. Additionally, establishing and implementing a dedicated requirements section would enhance functionality.
5As a provider or deployer of a high-risk AI system, I need to have a sub-service in the QMS that shows and stores the logs of the AI system’s use to comply with the regulations of the EU AIA and to ensure that no decisions are made with the AI system in critical situations.” (LR05). This requirement and the corresponding sub-service still need to be implemented.
6As a high-risk AI or GPAI system provider, I want to better understand how the created AI system works by applying explainability metrics. This will help me improve the system in future updates and create the best possible and most transparent instructions for end-users. These instructions shall inform them about good practices, limitations, and risks as comprehensively as possible.” (LR06). One explainability metric optimized for generative language models is implemented, helping users understand how much each input token influences the model’s generated output. This is useful for analyzing which types of tokens and grammar styles work better to achieve the best possible outcome. In future work, more types of explainability metrics will be added, not only for generative models but also for classification and regression models, allowing users of the QMS to select the most suitable metrics. Additionally, more visualizations will be researched and developed to better interpret the explainability metrics. Instructions on how to interpret these results will be provided through the QMS UI. However, no considerations have been made yet on how to automatically create or collaborate with humans to generate instructions for using the AI system. Such a feature will also be integrated into future versions of the QMS.
7As a user of the high-risk AI system, I want to have a UI to document risks and misuses and the potential errors and limitations of the AI system. In emergency cases, I want to be able to shut down the system.” (LR07). This is not yet implemented. The aim is to have the QMS available not only for providers of AI systems and deployers who modify the system and must comply with all regulations for high-risk AI systems but also the end-users of these AI systems. End-users can document limitations and risks within the QMS, which will be directly sent to the provider. The provider can then add this feedback to their backlog and address the issues in future updates. The QMS will be one extensive system with different functionalities, depending on the license, for example, provider or end-user.
8As an AI Expert, I want to evaluate the AI system, specifically the high-risk AI or GPAI system, on performance, robustness, explainability, security, and fairness to verify its use according to the EU AIA regulations and to identify its limitations.” (LR08). Five technical evaluation metrics for performance, explainability, and consistency have been implemented, allowing the user to evaluate the LLM in a specific domain for specific tasks. More metrics will be implemented, including those for classification and regression models. The design of the risk analysis component follows an object-oriented approach using the strategy software design pattern (cf. [6]). Additionally, no cybersecurity metrics have been implemented yet, as cybersecurity is only relevant for systems in production. However, integrating cybersecurity metrics will be addressed in future versions.
9As a provider or deployer of a high-risk AI or GPAI system, I want to check potential risks and perform model evaluations even aftermarket release to comply with the EU AIA and to detect potential new risks that have not been documented in the development phase.” (LR10). This requirement is fully fulfilled due to the design choice of building the QMS as a web application and integrating the AI system into the QMS. The risk assessment, creation of technical documentation, and all other features provided by the QMS can be accessed and used anytime without additional expenses.
10As an AI expert, domain expert, legal expert, or any other person involved in the assessment and documentation process of the high-risk AI or GPAI system, I want to have an application and user interface to easily interact with my colleagues, the model, and the tasks such as risk management and data governance, to ensure a transparent and reliable compliance management process, reduce costs, improve communication through a platform, and to create the proof of conformity according to the EU AIA regulations.” (SDR01, SDR02, SDR03). The presented design of the QMS is based on a SaaS approach. The idea is to have one system connected to an AI system that contains multiple sub-services, ideally one for each EU AIA regulation. The results from each sub-service are stored persistently in a database and can be printed as PDF documentation. However, some limitations still exist, such as the absence of user roles and the lack of support for collaboration, communication, and interaction among multiple users within the same process, such as in a risk assessment scenario. These features will be implemented in the second version of the prototype QMS.
NFR Evaluation - Design, Memory, and Performance. The QMS utilizes a microservice architecture. Each sub-service can be adopted, extended, and maintained independently of the others, guaranteeing a loosely coupled, and modular design (SDR05). All sub-services communicate via REST through the API Gateway with the frontend. Additionally, each sub-service backend communicates with its database using CRUD operations, ensuring persistent storage of all user and QMS data. (SDR06 and SDR07). Analyzing LLMs as a type of AI system is a good choice for determining the required GPU VRAM (memory), GPU, and CPU performance because these AI systems (models) are significantly larger than other AI systems. If the GPU and CPU resources are sufficient for LLMs, they should be more than adequate for most other AI systems and models. GPU VRAM is crucial for computing the gradients of LLMs and obtaining results for the presented explainability and consistency metrics. This is because all gradients are calculated at once, meaning that all parameters must be stored in the GPU’s VRAM simultaneously. The following calculation determines the required GPU VRAM for inferencing, gradient calculations, and one optimization step using a batch size of one. For instance, a 7B model comprises seven billion parameters (each including a weight and a bias term). Calculating the logits and gradients with full precision, meaning in float32, requires that each parameter be represented by 32 bits or 4 bytes. Therefore, a single forward pass necessitates: $7*10^{9}\>parameters*4\>bytes=28*10^{9}\>bytes$ (= 28GB). The same amount is again required for gradient computation, meaning an additional 28 GB is needed. Thus, approximately 56 GB of GPU VRAM is required for inference and gradient calculation. Switching to float16 precision from float32 has a significant impact on memory requirements. With the memory requirement for each parameter halved, only 28GB of memory is needed to compute a forward step and the gradients for a batch size of one. This underscores the potential for memory optimization and its impact on the overall system performance. By implementing optimizations such as gradient accumulation and other techniques, memory usage can be further reduced. The prototype QMS uses an NVIDIA RTX 4090 GPU with 24GB of GPU VRAM. This amount is insufficient for 7B LLMs like Meta’s LLama2-7B, and all larger Meta LLMs such as LLama3-8B. Therefore, a smaller model, such as Microsoft’s Phi3-mini, OpenAI’s GPT-2, and GPT-neo are integrated into the QMS. Phi-3-mini, which stores between 4 and 5 billion parameters, performs well in test cases, providing good results for all five technical evaluation metrics in 10 to 20 seconds. The Phi3-mini requires at most 20GB of GPU VRAM, making it suitable for the NVIDIA RTX 4090. The smaller language models can also be used as needed but do not deliver good results in summarization tasks. (SDR08, SDR09). In future work, memory and performance optimization will be further explored. Additionally, tests with real-life data will be conducted to evaluate the effectiveness of the metrics and the performance of different LLMs. Furthermore, other types of AI systems will be integrated such as neural networks for classification tasks. These models are significantly smaller than LLMs and can therefore be well-evaluated using the NVIDIA RTX4090 GPU.

### v2

@@ 3.4 Prototype QMS - First Version
To sign in, the user must enter test as username and password. The homepage is subdivided and structured into the prototype’s several sub-services. Each functional sub-service (RMS and DMDGS) is listed in a separate row, one below the other, and contains several boxes each, one for performing risk assessments or data checks and one for viewing past risk assessments or data checks. Additionally, the UI provides access to various sections with details about the LLM and the AIA.
@@
Sub-Service 01: Risk Management System. As mentioned, the main focus of the prototype QMS was on the RMS sub-service. The overall structure of the RMS is based on the ISO 31000 standard [22], which is also applied to the AI system RMS presented in [49]. It follows a “Plan-Do-Act-Check” principle: Plan involves planning the RMS based on the specific use case and defining the processes’ sub-activities, Do encompasses conducting the designed RMS process, Act entails inspecting and evaluating the entire RMS process, and output, and Check involves refining and optimizing the RMS process. Additionally, [34] establishes five design principles for RMS for AI systems, which should also be integrated. These principles include i) incorporating multi-perspective expert assessment, such as drawing insights from various domains and involving AI experts in the process, ii) encouraging participation from diverse stakeholders in the risk assessment process, iii) identifying risks based on real-life scenarios, and iv) analyzing risks using metrics beyond accuracy, v) ensuring “human in the loop” processes for black-box models. This RMS can be seen as a prototype within the prototype QMS. The construction process of this RMS, according to ISO 31000, consists of an iterative process containing six sub-activities: component selection, risk identification, verification data selection, risk analysis, risk assessment, and risk mitigation. In the component selection, the users should be encouraged to select the analyzed AI system (in this case, the LLM) and the task the AI system performs. In the second step, the user performs risk identification based on the risk categorization strategy for AI systems presented by [15]. A vocabulary-based approach is used where the user adds values for the LLM’s domain, purpose, capabilities, LLM user, and LLM subject. Based on these selected values, the risk class of the AI system is determined. A basic algorithm was implemented to evaluate the risk class according to the categorization presented by [15]. Future work will explore changing the risk identification process to a more stakeholder-oriented approach, as mentioned in the design principle by [34]. Additionally, it will be considered whether risk identification should be designed to identify foreseeable risks and misuses, assuming the AI system is already classified as high-risk, instead of categorizing the AI system into a risk class. In the third step, the user can quantitatively analyze the AI system’s performance, explainability, and consistency risks. The user can choose between different technical evaluation metrics from the risk analysis to the selected LLM. In the fourth step, the model calculates a risk assessment documentation based on the risk identification and analysis results. This documentation aims to verify compliance with AIA regulations. The current documentation will not serve as proof. However, the basic principle and the feasibility of automatically creating and downloading a risk and technical documentation from the prototype QMS can be demonstrated. In the final step, users can add strategies to mitigate assessed risks. This feature must still be fully developed and integrated into the prototype’s second version.
@@
The process was demonstrated with test data to provide an example of how such a risk analysis and assessment result looks, and all presented metrics were selected to be computed. As LLM, the Phi-3-mini-128k-instruct from Microsoft containing 3.8 billion parameters was analyzed on the domain “Industry Process Description”, and the task “Summarization”. The model input was: “Extract the Actor and Activity pairs from the text. Return only the list of JSON documents in the following format: [’actor’: ’example_actor_1’, ’activity’: ’example_activity_1’, ’actor’: ’example_actor_2’, ’activity’: ’example_activity_2’, …] without any further explanation: The user creates a new process instance, then the system can execute the instance and stop it afterward.”.
In Fig. 5, the performance results of the model from the risk assessment UI page are shown. The performance results display the numerical values for the model’s accuracy, Rouge score, and perplexity for the given input tasks.
@@
In Fig. 6, the explainability result is visualized as a saliency map, where each input token is colored between red (unimportant, less sensitive) and green (important, sensitive) to the model’s generated output.
@@
Sub-Service 02: Data Management and Governance System. The DMDGS is the second sub-service in the prototype QMS and is kept basic. It consists of two components in the UI: the data check page and the past data check page. The former allows the user to select the AI system (in this case, LLM) for which training, validation, or testing data has been used. Furthermore, the user can specify the name of the dataset, the origin, the type, the domain, and the size. To prove that the data is checked and compliant under Art. 10 AIA, the user has a text field in which he/ she can enter a textual reference to the proof. In future work, the goal is to enhance the Data Management and Governance System (DMDGS) to provide functionality for analyzing and assessing datasets for training, validation, and quantitative testing within the QMS. This includes evaluating datasets for biases and errors. Additionally, the QMS should verify that dataset splits are well-determined and that the data aligns with the goals of AI system development. It is also essential to check whether other EU data regulations, such as the GDPR, must be considered within the QMS.
@@ 4 Evaluation
First, the requirements and implementation of the initial prototype are evaluated. Next, the design of the prototype is assessed through a workshop with IT and AI experts and an interview with a legal expert.
4.1 Requirements
The proof-of-concept prototype demonstrates the feasibility of the proposed QMS design concept that connects to an AI system for technical assessments and incorporates multiple sub-services, each serving a specific purpose. Human involvement is facilitated through a web-based, service-oriented application. The QMS prototype utilizes a microservice architecture, enabling a modular and loosely coupled design to integrate multiple sub-services. The first-version prototype QMS incorporates no standard protocol requirements derived from the ISO 4213 [24] for AI performance evaluation, ISO CD TS 6254 [25] for AI explainability evaluation, and ISO TR 24029 for AI robustness evaluation [23]. However, these might be relevant to be compliant with more technical AIA requirements such as Art. 15 AIA. The metrics and calculations are currently included in the risk analysis component. However, a different sub-service can also be created for LR08 (Article 15 EU AIA) to prove the AI system’s accuracy, robustness, and cybersecurity separately. Computational and memory demands pose challenges, particularly for huge AI models like LLMs. For instance, the Phi-3-mini-128k-instruct, with 3.8 billion parameters, requires $2\cdot 3.8\cdot 10^{9}\>parameters\cdot 4\>bytes=30.4\cdot 10^{9}\>bytes$ = 30.4 GB of memory for a single forward and backward pass at full float32 precision. Reducing precision to float16 halves this requirement. The NVIDIA RTX 4090 GPU, with 24 GB of memory, can accommodate this LLM, which requires 15.2 GB in float16 to compute the implemented robustness and explainability metrics. Future work will focus on developing comprehensive, complete, and verifiable sub-services. To achieve this, ontology-based, vectorized database schemes may be implemented as proposed by [19], enabling the detailed design of sub-services tailored to AIA requirements and enriched with relevant information from associated ISO standards. This approach will facilitate the implementation and execution of verified processes through workflow engines, as advocated by [36].
4.2 Workshop with IT and AI Experts
A workshop was conducted to evaluate the prototype following the guidelines and principles of [48].
Method. Seven experts in information systems, business process management, and AI participated in the workshop. Three of them are continually involved with AI, especially LLMs, in their research. First, a presentation and prototype QMS live demo were provided, continuing with a focus-group discussion. The goal of the workshop was to gather feedback on the first version of the prototype QMS and to identify additional requirements for the RMS or the related future sub-service that allows them to technically assess the AI systems such that the QMS can be used to verify, evaluate and document AI systems entirely on different tasks. Therefore, the focus was more on evaluating the integrated RMS, technical metrics, and assessment capabilities.
Result. Five of seven experts agreed that such a tool would be helpful and increase efficiency when technically evaluating AI systems on specialized tasks, even if they mentioned that complex tasks might require manual inspection. LLMs whose output cannot have a pre-defined ground truth could be incredibly challenging for the QMS. Two had no opinion on such a tool since they never researched or worked on AI and needed to be more familiar with the AIA. Additionally, helpful technical requirements are elicited to improve the prototype QMS’s usefulness: i) the prototype QMS users shall be able to upload and store their data to analyze and assess the risk of the AI system (cf. Art. 9, 15 AIA). The data referred to by the experts pertains solely to the data used for evaluating the AI system during the post-development & pre-release phase and is unrelated to the training data. ii) Comprehensive data buckets of similar data should be automatically constructed based on the type and domain of data QMS users add. iii) Complete coverage of test cases is essential, meaning that all possible scenarios and edge cases must be considered and tested to guarantee the reliability and effectiveness of the AI system. This implies that the data used for evaluation should be comprehensive, mostly error-free, and cover real-life settings. iv) Clarification shall be needed, noted, and provided to the QMS user regarding manual procedures that cannot be automated within the QMS.
4.3 Inteview with Legal Expert
To evaluate the prototype, an interview was conducted with a legal expert and a corresponding qualitative content analysis (QCA) following [14].
Method. The QCA consists of a structured process containing four main steps: First, a research question (pp. 61-70) and a theoretical preliminary idea are constructed containing hypotheses built on a search pattern (pp. 73-90). The search pattern contains independent (implementation and conformance with AIA regulations, technical design, and incorporation of other EU regulations), intermediate (Prototype QMS), and dependent (ease verification and documentation of AI, improve efficiency in AI Compliance Management) variables. Based on having causal mechanisms between independent, intermediate, and dependent variables, a hypothesis was constructed: The implementation of AIA regulations as software requirements and the architecture influence the design of the prototype QMS, which can increase the efficiency for companies in the AI compliance management process which is either proved or disproved as a result of the QCA. Second, the interview partners are acquired, and an interview guideline is set up (pp. 111-150). Third, expert interviews are transliterated, extracted, analyzed, and categorized into the search pattern categories (pp. 197-257). Fourth, the results are interpreted and presented (pp. 261- 275).
Result. The legal expert evaluated the prototype positively. It was mentioned that the QMS has a very good, precise design and is, therefore, easy to use, even for auditors with legal expertise but possibly less AI expertise, such as potential auditors for national authorities or internal compliance team members and auditors. Therefore, one legal expert confirmed the hypothesis that the design concept and the prototype QMS positively impact AI systems’ verification, documentation, and compliance management process. Furthermore, it was confirmed that the QMS should be broad enough to accommodate, audit, and document all types of AI systems, including those not classified as high-risk. This approach would simplify the internal compliance management process for any AI system. Additionally, helpful technical requirements are elicited to improve the prototype QMS usefulness: i) The QMS shall technically check and verify AI systems according to generalized standard protocols such as ISO standards. ii) The QMS should adapt to changes in AI systems by focusing checks solely on modifications, with the results integrated into the existing documentation. iii) Documentation (cf. Art. 11 AIA) and potentially the transparency obligations (cf. Art. 14 AIA) shall not be displayed as separate, independent sub-services. Instead, all other sub-service results shall be selectable and added to the technical documentation. The technical documentation shall be a collection of all sub-service results. iv) Alternating reactions between the AI system, the AIA, and other harmonization standards (cf. Annex I AIA) depending on the domain of the AI system could arise and need to be checked.
To further enhance the prototype QMS, future work will expand qualitative evaluations by engaging a larger group of IT, AI, and legal experts. This will build on the insights provided by evaluating the first version of the prototype QMS, which only involves seven IT and AI experts and one legal expert. The opinion of further legal experts would help to refine the requirements and usability of the QMS and manifest the constructed hypothesis.